# Quantifying Uncertainty: All We Need is the Bootstrap?

### v1

@@ 2.2 Measuring the quality of confidence intervals
Most related work measures only coverage, with only a few studies measuring interval length [7, 11, 18, 21, 26, 37, 39] or comparing CIs with exact intervals [12, 14, 18]. While we also focus on coverage, we measure and report some results for interval length for two-sided CIs and the absolute distance from exact intervals for one-sided intervals. The exact interval for endpoint $\alpha$ and parameter $\theta$ is defined as $\hat{\theta}_{exact}[\alpha]=\hat{\theta}-\hat{\sigma}K^{-1}(1-\alpha)$, where $K$ is the cumulative distribution function of $\theta$ [13]. We approximate $\hat{\sigma}K^{-1}(1-\alpha)$ with 100000 samples.
Note that two-sided coverage can be misleading regarding a method’s coverage, because good two-sided coverage can be, and in practice often is, a result of substantial but opposite errors in the two one-sided intervals (see DiCiccio and Efron [13] for an example). That is, while two-sided error can be studied from one-sided CIs, the converse is not true.
We argue that the practical meaning of coverage error depends on nominal coverage and is not symmetric. For example, 51% coverage at 50% nominal coverage is not the same as 96% coverage at 95% nominal coverage. And 85% coverage is not the same as 95% coverage at 90% nominal coverage. To aggregate results and for a threshold-based criterion that can be applied to all confidence levels, we use Kullback-Leibler divergence - the information loss if we assume nominal coverage $\pi$ when actual coverage is $p$:
@@
For a threshold-based criterion of what we consider good enough we modify Bradley’s criterion $|p-\pi|<\frac{\min(\pi,1-\pi)}{k}$, where $\pi$ is the nominal coverage [5]. Common choices for $k$ are 10 (stringent), 4 (intermediate), 2 (liberal), and 0.75 (very liberal). Note that the intermediate and very liberal were introduced by Robey and Barcikowski [41]). At $95\%$ nominal coverage, the KL divergences for the Bradley lower bounds $94.5\%$ ($k=10$), $93.8\%$ ($k=4$), $92.5\%$ ($k=2$), and $88.3\%$ ($k=0.75$) are approximately $0.0004$, $0.0020$, $0.0083$, and $0.0503$. We adopt $\text{KL}(0.945,0.95)$ as the stringent criterion and a factor of 5 as an order of magnitude worse/better performance. For nominal coverage 95% this leads to criteria very similar to Bradley’s: $(93.8,96.2)$ vs $(93.9,96.1)$ (intermediate), $(92.5,97.5)$ vs $(92.4,97.3)$ (liberal), and $(88.3,101.7)$ vs $(88.6,99.4)$ (very liberal). However, for nominal coverage further away from 95% our approach gives more sensible criteria and does not produce endpoints outside of the unit interval.
@@ 3 Results
The simulation study consists of 1302 combinations and it is infeasible to list them all. We will focus on identifying if a bootstrap method is a viable one-size-fits-all approach. Inevitably, some details that are relevant to the reader will be left out. We prepared a visualization tool to browse the results of the experiments zrimseku.github.io/bootstrap-simulation/.
The raw results (aggregated over 10000 replications) and the source code for the tool, simulation framework, pre-processing, and analysis can be found here: github.com/zrimseku/bootstrap-simulation. The library with all the bootstrap methods can be found here: github.com/zrimseku/bootstrap-ci.
@@ 3.2 Coverage of bootstrap methods
Table 2 shows a comparison of bootstrap methods in mean KL. As expected, coverage improves with sample size and the two extreme percentiles and standard deviation are the most difficult functionals. For Pearson correlation, mean, and median, DB is best. For the percentiles, B-n is best. And for standard deviation, B-t is best. B-n and B-t perform the best overall. Because coverage gets wors with smaller sample size, the overall results are biased towards methods that perform well on small $n$. DB and BCa, which we would expect to perform best, perform relatively poorly for small $n$, but are best and second best for sample sizes $n\geq 64$.
@@
Table 3: Threshold-based coverage performance of bootstrapping on one-sided CIs. The table shows the percentage of experiments where a method does not meet the stringent criterion $\text{KL}(0.945,0.95)$. The overall column is across all experiments, while the remaining results are grouped by sample size or statistical functional. The best performing method for each column is underlined.
overall 4 8 16 32 64 128 256 corr mean median Q(0.05) Q(0.95) std
DB $\underline{0.64}$ $0.96$ $\underline{0.85}$ $\underline{0.82}$ $\underline{0.73}$ $\underline{0.59}$ $\underline{0.48}$ $\underline{0.20}$ $\underline{0.30}$ $\underline{0.52}$ $0.38$ $0.87$ $\underline{0.78}$ $0.75$
BCa $0.78$ $0.99$ $0.91$ $0.89$ $0.81$ $0.72$ $0.64$ $0.59$ $0.43$ $0.60$ $0.85$ $0.97$ $0.88$ $\underline{0.69}$
B-n $0.78$ $0.97$ $0.93$ $0.89$ $0.82$ $0.71$ $0.67$ $0.59$ $0.67$ $0.76$ $0.65$ $\underline{0.71}$ $0.80$ $1.00$
SB $0.79$ $0.96$ $0.94$ $0.83$ $0.78$ $0.74$ $0.71$ $0.66$ $0.67$ $0.75$ $\underline{0.33}$ $0.98$ $0.97$ $1.00$
PB $0.82$ $0.94$ $0.97$ $0.94$ $0.84$ $0.73$ $0.72$ $0.68$ $0.70$ $0.76$ $0.46$ $0.97$ $0.98$ $1.00$
BC $0.84$ $0.99$ $0.94$ $0.94$ $0.91$ $0.83$ $0.75$ $0.61$ $0.43$ $0.67$ $0.85$ $0.86$ $1.00$ $0.89$
B-t $0.85$ $\underline{0.89}$ $0.88$ $0.87$ $0.87$ $0.84$ $0.84$ $0.81$ $0.80$ $0.64$ $0.91$ $0.92$ $0.95$ $0.88$
BB $0.88$ $0.93$ $0.96$ $0.92$ $0.89$ $0.83$ $0.84$ $0.80$ $0.77$ $0.72$ $0.93$ $0.94$ $0.97$ $0.86$
Table 3 shows how often methods fail to meet the stringent criterion. DB outperforms other methods overall and on all sample sizes except $n=4$. It is the best method or relatively close to the best on all functionals, except on one of the two extreme percentiles. BCa is second-best overall. The results so far suggest that overall DB is best, but can have very poor coverage in some cases, especially for the two extreme percentiles at small sample sizes. While a more liberal criterion will result in fewer failures for all methods, the ordering does not change.
From the six confidence levels in our experiments we derive results for the two most common two-sided CIs (95%, 90%). Table 4 shows how often the methods fail to meet the liberal criterion for these two-sided intervals. These and results for other thresholds and KL are similar for one-sided and two-sided intervals. Most bootstrap methods meet the liberal criterion in almost all cases when $n\geq 128$ ($n\geq 64$ for DB), but even at $n=256$ there are still two cases where even DB does not (standard deviation for log-normal). Excluding these, DB meets the very liberal criterion for all experiments for $n\geq 64$.
@@
Table 4: Threshold-based coverage performance of bootstrapping on two-sided CIs. The table shows the percentage of experiments where a method does not meet the liberal criterion $25\times\text{KL}(0.945,0.95)$. The overall column is across all experiments, while the remaining results are grouped by sample size or statistical functional. The best performing method for each column is underlined.
overall 4 8 16 32 64 128 256 corr mean median Q(0.05) Q(0.95) std
DB $\underline{0.33}$ $0.81$ $\underline{0.50}$ $\underline{0.58}$ $\underline{0.52}$ $\underline{0.10}$ $\underline{0.03}$ $\underline{0.03}$ $\underline{0.00}$ $0.25$ $\underline{0.07}$ $0.43$ $\underline{0.40}$ $\underline{0.57}$
SB $0.45$ $1.00$ $0.67$ $0.73$ $0.63$ $0.24$ $0.13$ $0.08$ $0.10$ $0.46$ $0.14$ $0.43$ $0.42$ $0.82$
B-n $0.46$ $0.97$ $0.81$ $0.76$ $0.56$ $0.29$ $0.13$ $0.08$ $0.40$ $0.48$ $0.23$ $\underline{0.35}$ $0.43$ $0.80$
BCa $0.48$ $1.00$ $0.89$ $0.69$ $0.63$ $0.35$ $0.11$ $0.06$ $\underline{0.00}$ $0.46$ $0.25$ $0.62$ $\underline{0.40}$ $0.74$
PB $0.48$ $1.00$ $0.78$ $0.77$ $0.63$ $0.32$ $0.13$ $0.08$ $0.20$ $0.49$ $0.19$ $0.47$ $0.47$ $0.82$
BC $0.49$ $1.00$ $0.92$ $0.74$ $0.61$ $0.39$ $0.10$ $0.06$ $\underline{0.00}$ $0.48$ $0.25$ $0.57$ $0.45$ $0.77$
B-t $0.67$ $\underline{0.50}$ $0.72$ $0.81$ $0.76$ $0.66$ $0.61$ $0.60$ $0.40$ $\underline{0.17}$ $0.76$ $1.00$ $1.00$ $0.65$
BB $0.82$ $1.00$ $1.00$ $0.94$ $0.85$ $0.79$ $0.68$ $0.63$ $0.60$ $0.49$ $0.95$ $1.00$ $1.00$ $0.79$
@@ 3.3 Coverage comparison with baseline methods
Table 5: Cases where DB and baseline outperform each other. The number in the parentheses is the number of combinations where a method outperforms. There are a total of 36 combinations for every pair of $n$ and functional, except Pearson correlation, where there are 6.
n functional baseline $\gg$ DB DB $\gg$ baseline
4 corr fisher (6)
4 mean t-test (8); wilcoxon (1) DB (4)
4 median m-j (8); q-par (3)
4 Q(0.05) q-nonpar (12); q-par (22)
4 Q(0.95) q-par (21) DB (4)
4 std chi-sq (14) DB (6)
8 corr fisher (6)
8 mean t-test (4); wilcoxon (6) DB (4)
8 median m-j (2); q-par (1); wilcoxon (3) DB (4)
8 Q(0.05) q-nonpar (12); q-par (22) DB (3)
8 Q(0.95) q-nonpar (12); q-par (21)
8 std chi-sq (8) DB (11)
16 mean t-test (8); wilcoxon (8) DB (13)
16 median wilcoxon (1) DB (7)
16 Q(0.05) q-nonpar (17); q-par (15)
16 Q(0.95) m-j (11); q-par (11) DB (5)
16 std chi-sq (2) DB (12)
32 mean t-test (3); wilcoxon (3) DB (14)
32 median m-j (1) DB (3)
32 Q(0.05) m-j (6); q-nonpar (17); q-par (5)
32 Q(0.95) m-j (18); q-par (6)
32 std chi-sq (2) DB (15)
64 mean DB (12)
64 median DB (1)
64 Q(0.05) m-j (6); q-nonpar (3)
64 Q(0.95) m-j (5) DB (11)
64 std chi-sq (1) DB (20)
128 mean DB (14)
128 Q(0.05) m-j (9) DB (1)
128 Q(0.95) m-j (5) DB (1)
128 std DB (28)
256 mean DB (10)
256 Q(0.05) DB (4)
256 Q(0.95) DB (10)
256 std DB (27)
Table 5 adds detail to the results from Section 3.2. For Pearson correlation DB does not produce CIs for $n=4$ and performs poorly for $n=8$. It performs poorly on the two extreme percentiles for $n\leq 32$ and in some cases on the other functionals, mostly for $n\leq 8$. For Q(0.05) and Q(0.95), we would do better by choosing q-par for the smaller sample sizes and m-j for sample size $n=32$. For the other functionals and $n\geq 8$ ($n\geq 16$ for Pearson correlation) there is no clear advantage of using baseline methods. Note that for Q(0.05) and Q(0.95) and $n=16,32$ B-n is as good as or better than baseline. However, for $n\leq 8$ even B-n is worse than the baseline, but less so than DB.

### v2

@@ 2.2 Measuring the quality of confidence intervals
Most related work measures only coverage, with only a few studies measuring interval length [2, 37, 25, 35, 36, 17, 6] or comparing CIs with exact intervals [22, 14, 25].
While we also focus on coverage, we measure and report results for interval length for two-sided CIs and the absolute distance from exact intervals for one-sided intervals. The exact interval for endpoint $\alpha$ and parameter $\theta$ is defined as $\hat{\theta}_{exact}[\alpha]=\hat{\theta}-\hat{\sigma}K^{-1}(1-\alpha)$, where $K$ is the cumulative distribution function of $\theta$ [46]. We approximate $\hat{\sigma}K^{-1}(1-\alpha)$ with 100000 samples.
Note that two-sided coverage can be misleading regarding a method’s coverage, because good two-sided coverage can be, and in practice often is, a result of substantial, but opposite errors in the two one-sided intervals (see [46] for an example). That is, while two-sided error can be studied from one-sided CIs, the converse is not true.
We argue that the practical meaning of coverage error depends on nominal coverage and is not symmetric. For example, 51% coverage at 50% nominal coverage is not the same as 96% coverage at 95% nominal coverage. And 85% coverage is not the same as 95% coverage at 90% nominal coverage. To aggregate results and for a threshold-based criterion that can be applied to all confidence levels, we propose a a novel criterion based on the Kullback-Leibler divergence. That is, we measure information loss if we assume nominal coverage $\pi$ when actual coverage is $p$:
@@
For a threshold-based criterion of what we consider good enough we modify Bradley’s criterion $|p-\pi|<\frac{\min(\pi,1-\pi)}{k}$, where $\pi$ is the nominal coverage [47]. Common choices for $k$ are 10 (stringent), 4 (intermediate), 2 (liberal), and 0.75 (very liberal). Note that the intermediate and very liberal were introduced by [48]. In related work on the bootstrap, only [37] uses a good enough criterion - that actual coverage should lie between 92.5% and 97.5% when nominal coverage is 95% (this is based on the work of [47]).
At $95\%$ nominal coverage, the KL divergences for the Bradley lower bounds $94.5\%$ ($k=10$), $93.8\%$ ($k=4$), $92.5\%$ ($k=2$), and $88.3\%$ ($k=0.75$) are approximately $0.0004$, $0.0020$, $0.0083$, and $0.0503$. We adopt $\text{KL}(0.945,0.95)$ as the stringent criterion and a factor of 5 as an order of magnitude worse/better performance. For nominal coverage 95% this leads to criteria very similar to Bradley’s: $(93.8,96.2)$ vs $(93.9,96.1)$ (intermediate), $(92.5,97.5)$ vs $(92.4,97.3)$ (liberal), and $(88.3,101.7)$ vs $(88.6,99.4)$ (very liberal). However, for nominal coverage further away from 95% our approach gives produces sensible criteria and does not produce endpoints outside of the unit interval.
@@ 3 Results
The simulation study consists of 1302 combinations and it is infeasible to list them all. We will focus on identifying if a bootstrap method is a viable one-size-fits-all approach. Inevitably, some details that are relevant to the reader might be left out. We prepared a visualization tool to browse all of the the results of the experiments zrimseku.github.io/bootstrap-simulation/ (see Figure 2).
The raw results (aggregated over 10000 replications) and the source code for the tool, simulation framework, pre-processing, and analysis can be found here: github.com/zrimseku/bootstrap-simulation. The source code can be used to generate the full non-aggregated results. The library with all the bootstrap methods can be found here: github.com/zrimseku/bootstrap-ci.
@@ 3.2 Coverage of bootstrap methods
Table 3.2 shows a comparison of bootstrap methods in mean KL. As expected, coverage improves with sample size and the two extreme percentiles and standard deviation are the most difficult functionals. For Pearson correlation, mean, and median, DB is best. For the percentiles, B-n is best. And for standard deviation, B-t is best. B-n and B-t perform the best overall. Because coverage gets wors with smaller sample size, the overall results are biased towards methods that perform well on small $n$. DB and BCa, which we would expect to perform best, perform relatively poorly for small $n$, but are best and second best for sample sizes $n\geq 64$.
[…]
@@
Mean KL coverage performance of bootstrap methods for one-sided CIs. The all column is across all combinations, while the remaining results are grouped by sample size or statistical functional. The best performing method for each column is underlined. all 4 8 16 32 64 128 256 corr mean Q0.5 Q0.05 Q0.95 std B-n $\underline{0.082}$ $0.322$ $\underline{0.180}$ $\underline{0.091}$ $\underline{0.038}$ $0.021$ $0.014$ $0.010$ $0.006$ $0.030$ $0.004$ $\underline{0.024}$ $\underline{0.145}$ $0.208$ B-t $0.088$ $\underline{0.199}$ $0.280$ $0.112$ $0.051$ $0.024$ $0.020$ $0.014$ $0.008$ $0.015$ $0.012$ $0.078$ $0.264$ $\underline{0.084}$ $B$ $0.117$ $0.253$ $0.194$ $0.227$ $0.106$ $0.047$ $0.034$ $0.023$ $0.017$ $0.034$ $0.045$ $0.159$ $0.234$ $0.145$ SB $0.124$ $0.459$ $0.278$ $0.160$ $0.063$ $0.028$ $0.018$ $0.011$ $0.002$ $0.026$ $0.003$ $0.064$ $0.230$ $0.308$ DB $0.140$ $0.544$ $0.348$ $0.200$ $0.060$ $\underline{0.008}$ $\underline{0.005}$ $\underline{0.003}$ $\underline{0.000}$ $\underline{0.010}$ $\underline{0.002}$ $0.125$ $0.401$ $0.188$ BCa $0.159$ $0.600$ $0.335$ $0.242$ $0.096$ $0.018$ $0.010$ $0.005$ $0.001$ $0.023$ $0.017$ $0.201$ $0.379$ $0.218$ BC $0.162$ $0.607$ $0.350$ $0.234$ $0.096$ $0.022$ $0.013$ $0.008$ $0.001$ $0.026$ $0.017$ $0.169$ $0.387$ $0.246$ PB $0.165$ $0.620$ $0.378$ $0.229$ $0.076$ $0.029$ $0.019$ $0.012$ $0.002$ $0.028$ $0.003$ $0.109$ $0.377$ $0.330$
[…]
@@
Threshold-based coverage performance of bootstrapping on one-sided CIs. The table shows the percentage of experiments where a method does not meet the stringent criterion $\text{KL}(0.945,0.95)$. The all column is across all experiments, while the remaining results are grouped by sample size or statistical functional. The best performing method for each column is underlined. all 4 8 16 32 64 128 256 corr mean Q0.5 Q0.05 Q0.95 std DB $\underline{0.64}$ $0.96$ $\underline{0.85}$ $\underline{0.82}$ $\underline{0.73}$ $\underline{0.59}$ $\underline{0.48}$ $\underline{0.20}$ $\underline{0.30}$ $\underline{0.52}$ $0.38$ $0.87$ $\underline{0.78}$ $0.75$ BCa $0.78$ $0.99$ $0.91$ $0.89$ $0.81$ $0.72$ $0.64$ $0.59$ $0.43$ $0.60$ $0.85$ $0.97$ $0.88$ $\underline{0.69}$ B-n $0.78$ $0.97$ $0.93$ $0.89$ $0.82$ $0.71$ $0.67$ $0.59$ $0.67$ $0.76$ $0.65$ $\underline{0.71}$ $0.80$ $1.00$ SB $0.79$ $0.96$ $0.94$ $0.83$ $0.78$ $0.74$ $0.71$ $0.66$ $0.67$ $0.75$ $\underline{0.33}$ $0.98$ $0.97$ $1.00$ PB $0.82$ $0.94$ $0.97$ $0.94$ $0.84$ $0.73$ $0.72$ $0.68$ $0.70$ $0.76$ $0.46$ $0.97$ $0.98$ $1.00$ BC $0.84$ $0.99$ $0.94$ $0.94$ $0.91$ $0.83$ $0.75$ $0.61$ $0.43$ $0.67$ $0.85$ $0.86$ $1.00$ $0.89$ B-t $0.85$ $\underline{0.89}$ $0.88$ $0.87$ $0.87$ $0.84$ $0.84$ $0.81$ $0.80$ $0.64$ $0.91$ $0.92$ $0.95$ $0.88$ $B$ $0.88$ $0.93$ $0.96$ $0.92$ $0.89$ $0.83$ $0.84$ $0.80$ $0.77$ $0.72$ $0.93$ $0.94$ $0.97$ $0.86$
[…]
Threshold-based coverage performance of bootstrapping on two-sided CIs. The table shows the percentage of experiments where a method does not meet the liberal criterion $25\times\text{KL}(0.945,0.95)$. The all column is across all experiments, while the remaining results are grouped by sample size or statistical functional. The best performing method for each column is underlined. all 4 8 16 32 64 128 256 corr mean Q0.5 Q0.05 Q0.95 std DB $\underline{0.33}$ $0.81$ $\underline{0.50}$ $\underline{0.58}$ $\underline{0.52}$ $\underline{0.10}$ $\underline{0.03}$ $\underline{0.03}$ $\underline{0.00}$ $0.25$ $\underline{0.07}$ $0.43$ $\underline{0.40}$ $\underline{0.57}$ SB $0.45$ $1.00$ $0.67$ $0.73$ $0.63$ $0.24$ $0.13$ $0.08$ $0.10$ $0.46$ $0.14$ $0.43$ $0.42$ $0.82$ B-n $0.46$ $0.97$ $0.81$ $0.76$ $0.56$ $0.29$ $0.13$ $0.08$ $0.40$ $0.48$ $0.23$ $\underline{0.35}$ $0.43$ $0.80$ BCa $0.48$ $1.00$ $0.89$ $0.69$ $0.63$ $0.35$ $0.11$ $0.06$ $\underline{0.00}$ $0.46$ $0.25$ $0.62$ $\underline{0.40}$ $0.74$ PB $0.48$ $1.00$ $0.78$ $0.77$ $0.63$ $0.32$ $0.13$ $0.08$ $0.20$ $0.49$ $0.19$ $0.47$ $0.47$ $0.82$ BC $0.49$ $1.00$ $0.92$ $0.74$ $0.61$ $0.39$ $0.10$ $0.06$ $\underline{0.00}$ $0.48$ $0.25$ $0.57$ $0.45$ $0.77$ B-t $0.67$ $\underline{0.50}$ $0.72$ $0.81$ $0.76$ $0.66$ $0.61$ $0.60$ $0.40$ $\underline{0.17}$ $0.76$ $1.00$ $1.00$ $0.65$ $B$ $0.82$ $1.00$ $1.00$ $0.94$ $0.85$ $0.79$ $0.68$ $0.63$ $0.60$ $0.49$ $0.95$ $1.00$ $1.00$ $0.79$
Table 3.2 shows how often methods fail to meet the stringent criterion. DB outperforms other methods overall and on all sample sizes except $n=4$. It is the best method or relatively close to the best on all functionals, except on one of the two extreme percentiles. BCa is second-best overall. The results so far suggest that overall DB is best, but can have very poor coverage in some cases, especially for the two extreme percentiles at small sample sizes. While a more liberal criterion will result in fewer failures for all methods, the ordering does not change.
From the six confidence levels in our experiments we derive results for the two most common two-sided CIs (95%, 90%). Table 3.2 shows how often the methods fail to meet the liberal criterion for these two-sided intervals. These and results for other thresholds and KL are similar for one-sided and two-sided intervals. Most bootstrap methods meet the liberal criterion in almost all cases when $n\geq 128$ ($n\geq 64$ for DB), but even at $n=256$ there are still two cases where even DB does not (standard deviation for log-normal). Excluding these, DB meets the very liberal criterion for all experiments for $n\geq 64$.
@@ 3.3 Coverage comparison with baseline methods
[…]
Cases where DB and baseline outperform each other. The number in the parentheses is the number of combinations where a method outperforms. There are a total of 36 combinations for every pair of $n$ and functional, except Pearson correlation, where there are 6. n functional baseline $\gg$ DB DB $\gg$ baseline 4 corr fisher (6) 4 mean t-test (8); wilcoxon (1) DB (4) 4 median m-j (8); q-par (3) 4 Q(0.05) q-nonpar (12); q-par (22) 4 Q(0.95) q-par (21) DB (4) 4 std chi-sq (14) DB (6) 8 corr fisher (6) 8 mean t-test (4); wilcoxon (6) DB (4) 8 median m-j (2); q-par (1); wilcoxon (3) DB (4) 8 Q(0.05) q-nonpar (12); q-par (22) DB (3) 8 Q(0.95) q-nonpar (12); q-par (21) 8 std chi-sq (8) DB (11) 16 mean t-test (8); wilcoxon (8) DB (13) 16 median wilcoxon (1) DB (7) 16 Q(0.05) q-nonpar (17); q-par (15) 16 Q(0.95) m-j (11); q-par (11) DB (5) 16 std chi-sq (2) DB (12) 32 mean t-test (3); wilcoxon (3) DB (14) 32 median m-j (1) DB (3) 32 Q(0.05) m-j (6); q-nonpar (17); q-par (5) 32 Q(0.95) m-j (18); q-par (6) 32 std chi-sq (2) DB (15) 64 mean DB (12) 64 median DB (1) 64 Q(0.05) m-j (6); q-nonpar (3) 64 Q(0.95) m-j (5) DB (11) 64 std chi-sq (1) DB (20) 128 mean DB (14) 128 Q(0.05) m-j (9) DB (1) 128 Q(0.95) m-j (5) DB (1) 128 std DB (28) 256 mean DB (10) 256 Q(0.05) DB (4) 256 Q(0.95) DB (10) 256 std DB (27)
Table 3.3 adds detail to the results from Section 3.2. For Pearson correlation DB does not produce CIs for $n=4$ and performs poorly for $n=8$. It performs poorly on the two extreme percentiles for $n\leq 32$ and in some cases on the other functionals, mostly for $n\leq 8$. For Q(0.05) and Q(0.95), we would do better by choosing q-par for the smaller sample sizes and m-j for sample size $n=32$. For the other functionals and $n\geq 8$ ($n\geq 16$ for Pearson correlation) there is no clear advantage of using baseline methods. Note that for Q(0.05) and Q(0.95) and $n=16,32$ B-n is as good as or better than baseline. However, for $n\leq 8$ even B-n is worse than the baseline, but less so than DB.
Table 3.3 adds detail to the results from Section 3.2. For Pearson correlation DB does not produce CIs for $n=4$ and performs poorly for $n=8$. It performs poorly on the two extreme percentiles for $n\leq 32$ and in some cases on the other functionals, mostly for $n\leq 8$. For Q(0.05) and Q(0.95), we would do better by choosing q-par for the smaller sample sizes and m-j for sample size $n=32$. For the other functionals and $n\geq 8$ ($n\geq 16$ for Pearson correlation) there is no clear advantage of using baseline methods. Table 3.3 shows the advantage of using B-n instead of DB for Q(0.05) and Q(0.95) for $n\leq 32$.
[…]
Cases where B-n and baseline outperform each other. Results are for the two extreme percentiles and $n\leq 32$. The number in the parentheses is the number of combinations where a method outperforms. There are a total of 36 combinations for every pair of $n$ and functional. n functional baseline $\gg$ B-n B-n $\gg$ baseline 4 Q(0.05) q-nonpar (11); q-par (14) 4 Q(0.95) q-par (14) 8 Q(0.05) q-par (13) B-n (1) 8 Q(0.95) q-nonpar (4); q-par (13) B-n (4) 16 Q(0.05) q-nonpar (5); q-par (6) B-n (5) 16 Q(0.95) m-j (2); q-par (6) B-n (2) 32 Q(0.05) B-n (10) 32 Q(0.95) m-j (3); q-par (1) B-n (11)